OpenStudy (anonymous):

A jet flies 3,360 miles from New York to Paris in about 2.7 hours. About how many hours would the jet take to fly 5,687 miles from Paris to Rio de Janeiro?

OpenStudy (a1234):

Proportion: 2.7/3360 = x/5687 15354.9/3360 = Approximately 4.6 hours.